# FuseMoE: Mixture-of-Experts Transformers for Fleximodal Fusion

### v2

@@ 4.2 Ablation Studies
Figure 4(c) illustrates the effectiveness of utilizing per-modality routers and the entropy loss $\mathcal{E}$ in addressing missing modalities. Initially, we compare the performance of FuseMoE on patients with fully available modalities against those with missing components, employing a joint router mechanism with the importance loss function (shazeer2017outrageously, ), to ensure load balancing. The inclusion of datasets with missing modalities, while expanding the sample size, resulted in a decrease in performance due to the compromised data quality. However, a performance enhancement was observed upon integrating per-modality or disjoint routers with $\mathcal{E}$. Notably, the outcomes for the 48-IHM and LOS tasks with missing modalities surpassed those obtained from datasets without any missingness. This is because the per-modality approach can better separate the present and missing modalities, reducing the influence of experts responsible for processing the absent inputs. Therefore, this leads to a more efficient exploitation of a broader array of samples.
@@ Appendix A Related Works
Initial approaches to multimodal fusion incorporated techniques such as kernel-based methods [7, 11, 68], graphical models [59, 21, 74], and neural networks [60, 20, 66]. With the diverse evolution of deep learning models, numerous advanced methods have now been employed in the fusion of multimodal data. In the realm of sentiment analysis, [94, 52] employ a low-rank Tensor Fusion method that leverages both language and video content. Attention-gating mechanisms are used by [72, 93] to generate displacement vectors through cross-modal self-attention, which are then added to the input vectors from the primary modality. [88] takes an alternative approach by integrating multiple layers of cross-modal attention blocks in a word-level vision/language/audio alignment task.
In the context of clinical prediction, [45, 16] adopt a late fusion approach to combining vital sign and text data by concatenating embeddings from pre-trained feature extractors. [81] developed a generalizable data preprocessing and modeling pipeline for EHR encompassing four data modalities, albeit through a direct concatenation of existing feature embeddings for each modality followed by an XGBoost classifier [12]. Recently, [99] expanded on the work of [88] by introducing a discretized multi-time attention (mTAND) module [80] to encode temporal irregularities in time series and text data. Their fusion approach involves layering sets of self- and cross-modal attention blocks. However, this approach is limited to just two modalities and is not easily extendable to include additional modal components or handle missing modalities. To the best of our knowledge, existing works are tailored to application-specific settings that necessitate the computation of pairwise cross-modal relationships, which are not scalable to more general settings with arbitrary modalities. Moreover, these studies typically do not account for scenarios where modalities are missing, or rely on imputation approaches based on observed data.
@@
MoE [35, 91] has gained significant popularity for managing complex tasks since its introduction three decades ago. Unlike traditional models that reuse the same parameters for all inputs, MoE selects distinct parameters for each specific input. This results in a sparsely activated layer, enabling a substantial scaling of model capacity without a corresponding increase in computational cost. Recent studies have demonstrated the effectiveness of integrating MoE with cutting-edge models across a diverse range of tasks [77, 19, 101]. These works have also tackled key challenges such as accuracy and training instability [65, 102, 70]. Given its ability to assign input partitions to specialized experts, MoE naturally lends itself to multimodal applications. This approach has been explored in fields such as vision-language modeling [58, 78] and dynamic image fusion [9]. However, the application of MoE in complex real-world settings, such as those involving FlexiModal Data, remains largely unexplored. This gap presents an opportunity to leverage MoE’s potential in handling its intricate and multifaceted nature such as multimodal EHR, where reliable multimodal integration is crucial.
@@
While MoE has been widely employed to scale up large models, its theoretical foundations have remained nascent. Recently, [64] provided convergence rates for both density and parameter estimation of Softmax gating Gaussian MoE. They connected these rates to the solvability of systems of polynomial equations under Voronoi-based loss functions. Later, [63] extended these theories to top-K sparse softmax gating MoE. Their theories further characterize the effect of the sparsity of gating functions on the behaviors of parameter estimation and verify the benefits of using top-1 sparse softmax gating MoE in practice. Other theoretical results for MoE include estimation rates of parameters and experts for multinomial logistic MoE [62], for dense-to-sparse gating MoE [61], and for input-independent gating MoE [29].
@@ B.1 MIMIC-IV and MIMIC-III Datasets
In the ICU, where rapid and informed decisions are crucial, accurate mortality prediction is essential to provide clinicians with advanced warnings of patient deterioration, aiding in critical decision-making processes [5]. Similarly, the prediction of patient length-of-stay is indispensable for optimizing treatment plans, resource allocation, and discharge processes [6]. Further, phenotyping of critical care conditions is highly relevant to comorbidity detection and risk adjustment and presents a more challenging task than binary classification, due to the heterogeneous presentation of conditions and the larger number of prediction tasks [99]. We concentrate on three critical care tasks as highlighted in [44], performing extensive empirical analysis on each building block of the proposed framework.
@@
  • 25-PHE  In this multilabel classification problem, we attempt to predict one of 25 acute care conditions [18, 53] (e.g., congestive heart failure, pneumonia, shock, etc.) at the end each each patient’s ICU stay. Because the original task was designed for diagnoses based on ICD-9 codes, but MIMIC-IV includes both ICD-9 and ICD-10 codes, we map patients with diagnoses coded using ICD-10 using the conversion database provided by [8].
@@
We implement an in-hospital mortality prediction (48-IHM) task to evaluate our method’s ability to predict short-term patient deterioration. Similarly, an accurate determination of patient discharge times is crucial for optimizing patient outcomes and hospital resource allocation [6], which motivates our length-of-stay (LOS) task. We frame 48-IHM and LOS as binary classification problems and use a 48-hour observation window (for patients who spent at least 48 hours in the ICU) to predict in-hospital mortality (48-IHM) and discharge (without expiration) within the 48 hours following the observation window (LOS). Lastly, identifying the presence of specific acute care conditions in patient records is essential for various clinical objectives, including the construction of cohorts for clinical studies and the detection of comorbidities [1]. Traditional methods, often reliant on manual chart reviews or simple billing code-based definitions, are increasingly being supplemented by machine learning techniques [27]; automating this process requires high-fidelity classifications, motivating our 25-type phenotype classification (25-PHE) task. In this multilabel classification problem, we attempt to predict one of 25 acute care conditions using data from the entire ICU stay.
@@
In our initial analysis, we focused on patients with no missing modalities, resulting in a dataset comprised of 8,770 ICU stays for the 48-IHM and LOS tasks, and 14,541 stays for the 25-PHE task. For our analyses with missing observations, we include a total of 35,129 stays for 48-IHM and LOS, and 71.173 for 25-PHE. To evaluate the single-label tasks, 48-IHM and LOS, we employ the F1-score and AUROC as our primary metrics. In line with previous studies [99, 50, 3], we use macro-averaged F1-score and AUROC to assess the 25-PHE task.
@@
We leveraged data from MIMIC-IV [36], a comprehensive database with records from nearly $300k$ patients admitted to a medical center from 2008 to 2019, focusing on the subset of 73,181 ICU stays. We were able to link core ICU records (containing lab results and vital signs) to corresponding chest X-rays [39], radiological notes [38], and electrocardiogram (ECG) data [24] taking place during a given ICU stay. We allocated 70 percent of the data for model training, with the remaining 30 percent evenly split between validation and testing.
@@ B.2 MOSI and MOSEI Datasets
Following previous work such as [26], we adopt mean absolute error (MAE), Pearson correlation (Corr), binary classification accuracy, F1 score computed for non-negative/negative class as evaluation metrics.
@@ B.4 CIFAR-10 Dataset
CIFAR-10 [46] is an established computer-vision dataset used for object recognition. It consists of 60,000 32x32 color images containing one of 10 object classes (”plane”, ”car”, ”bird”, ”cat”, ”deer”, ”dog”, ”frog”, ”horse”, ”ship”, ”truck”), with 6000 images per class.
@@ D.1 MIMIC-IV
In the preprocessing stage, we focused on 30 pertinent lab and chart events from each patient’s ICU record for vital sign measurements. For chest X-rays, we utilized a pre-trained DenseNet-121 model [14], which was fine-tuned on the CheXpert dataset [34], to extract 1024-dimensional image embeddings. For radiological notes, we obtained 768-dimensional embeddings using the BioClinicalBERT model [2]. ECG signals were processed using a convolutional autoencoder, adapted from [4], to generate a 256-dimensional embedding for each ECG.
@@
We selected 30 time series events for inclusion, following [81]. Nine of these were vital signs: heart rate, mean/systolic/diastolic blood pressure, respiratory rate, oxygen saturation, and Glascow Coma Scale (GCS) verbal, eye, and motor response. We also included 21 lab values: potassium, sodium, chloride, creatinine, urea nitrogram, bicarbonate, anion gap, hemoglobin, hematocrit, magnesium, platlet count, phosphate, white blood cell count, total calcium, MCH, red blood cell count, MCHC, MCV, RDW, platlet count, neutrophil count, and vancomycin. We standard scale each time series value to have mean 00 and standard deviation $1$, based on the values in the training set.
@@
To incorporate a medical imaging modality into our analyses, we use the MIMIC-CXR-JPG [37] module available from Physionet [23], which includes 377,110 JPG format images derived from the DICOM-based MIMIC-CXR database [39]. Following [81], for each image, we resize each JPG image to 224 $\times$ 224 pixels and then extract embeddings from the last layer of the Densenet121 model. We identify X-rays taken while the patient was in the ICU by first matching subject IDs in MIMIC-CXR-JPG with the core MIMIC-IV database, then limiting these matched X-rays to those with a chart time occuring between an ICU admission and discharge.
@@
To incorporate text data, we use the MIMIC-IV-Note module [38], which contains 2,321,355 deidentified radiology reports for 237,427 patients that can be matched with patients in the main MIMIC-IV via a similar approach to chest X-rays. We note that we were unable to obtain intermediate clinical notes (i.e., notes made by clinicians throughout a patient stay), as those have not yet been publicly released. We extract note embeddings using Bio-Clinical BERT [2].
@@
To include ECGs as an additional modality in our models, we utilize the MIMIC-IV-ECG [24] module, which includes approximately 800,000 ECGs (10 seconds, sampled at 500 Hz) collected from nearly 160,000 unique patients. To transform the ECGs so that they are suitable for input to our model, we adopt a convolutional autoencoder approach, adapted from [4], that compresses each ECG into a 256-dimensional vector. Specifically, each diagnostic ECG contains a $5000\times 12$ dimensional vector (5000 time points $\times$ 12 ECG leads). To prepare the ECG for input to the autoencoder, we only include the first $4096$ time points. We then train the autoencoder to compress the ECG into a 256-dimensional latent vector, and then reconstruct the original ECG using upsampling layers, using mean squared error as our loss function. The architecture is shown in Figure 6. We train the autoencoder with 90% of the ECGs available in the MIMIC-IV-ECG projection and use the rest for validation. We selected a batch size of 2048, and reduced the learning rate by a factor of $0.5$ if the validation loss had plateaued for $3$ epochs. Training stopped if the validation loss had not decreased for $6$ epochs. For our encoder, we use filter numbers of $[16,16,32,32,64,64]$, kernel widths of $[5,5,5,3,3,3]$ and a dropout rate of $0.1$. For the decoder, we use the same filter numbers and kernel widths in reverse, and maintain a dropout rate of $0.1$.

### v4

@@ 4.2 Ablation Studies
Figure 4(c) illustrates the effectiveness of utilizing per-modality routers and the entropy loss $\mathcal{E}$ in addressing missing modalities. Initially, we compare the performance of FuseMoE on patients with fully available modalities against those with missing components, employing a joint router mechanism with the importance loss function (Shazeer et al., 2017), to ensure load balancing. The inclusion of datasets with missing modalities, while expanding the sample size, resulted in a decrease in performance due to the compromised data quality. However, a performance enhancement was observed upon integrating per-modality or disjoint routers with $\mathcal{E}$. Notably, the outcomes for the 48-IHM and LOS tasks with missing modalities surpassed those obtained from datasets without any missingness. This is because the per-modality approach can better separate the present and missing modalities, reducing the influence of experts responsible for processing the absent inputs. Therefore, this leads to a more efficient exploitation of a broader array of samples. Implementation of FuseMoE and our baselines can be found at https://github.com/aaronhan223/FuseMoE.
@@ 5 Discussions and Limitations
Acknowledgement
Xing Han and Suchi Saria acknowledge support from the National Science Foundation (NSF) and the Gordon and Betty Moore Foundation. Carl Harris acknowledges support from the NSF Graduate Research Fellowship under Grant No. DGE 2139757. Nhat Ho acknowledges support from the NSF IFML 2019844 and the NSF AI Institute for Foundations of Machine Learning. Any opinion, findings, and conclusions or recommendations expressed in this material are those of the authors(s) and do not necessarily reflect the views of the NSF, Gordon and Betty Moore Foundation.
@@ Appendix A Related Works
Initial approaches to multimodal fusion incorporated techniques such as kernel-based methods [Bucak et al., 2013, Chen et al., 2014, Poria et al., 2015], graphical models [Nefian et al., 2002, Garg et al., 2003, Reiter et al., 2007], and neural networks [Ngiam et al., 2011, Gao et al., 2015, Nojavanasghari et al., 2016]. With the diverse evolution of deep learning models, numerous advanced methods have now been employed in the fusion of multimodal data. In the realm of sentiment analysis, Zadeh et al. [2017], Liu et al. [2018] employ a low-rank Tensor Fusion method that leverages both language and video content. Attention-gating mechanisms are used by Rahman et al. [2020], Yang and Wu [2021] to generate displacement vectors through cross-modal self-attention, which are then added to the input vectors from the primary modality. Tsai et al. [2019] takes an alternative approach by integrating multiple layers of cross-modal attention blocks in a word-level vision/language/audio alignment task.
In the context of clinical prediction, Khadanga et al. [2019], Deznabi et al. [2021] adopt a late fusion approach to combining vital sign and text data by concatenating embeddings from pre-trained feature extractors. Soenksen et al. [2022] developed a generalizable data preprocessing and modeling pipeline for EHR encompassing four data modalities, albeit through a direct concatenation of existing feature embeddings for each modality followed by an XGBoost classifier Chen and Guestrin [2016]. Recently, Zhang et al. [2023] expanded on the work of Tsai et al. [2019] by introducing a discretized multi-time attention (mTAND) module Shukla and Marlin [2021] to encode temporal irregularities in time series and text data. Their fusion approach involves layering sets of self- and cross-modal attention blocks. However, this approach is limited to just two modalities and is not easily extendable to include additional modal components or handle missing modalities. To the best of our knowledge, existing works are tailored to application-specific settings that necessitate the computation of pairwise cross-modal relationships, which are not scalable to more general settings with arbitrary modalities. Moreover, these studies typically do not account for scenarios where modalities are missing, or rely on imputation approaches based on observed data.
@@
MoE [Jacobs et al., 1991, Xu et al., 1994] has gained significant popularity for managing complex tasks since its introduction three decades ago. Unlike traditional models that reuse the same parameters for all inputs, MoE selects distinct parameters for each specific input. This results in a sparsely activated layer, enabling a substantial scaling of model capacity without a corresponding increase in computational cost. Recent studies have demonstrated the effectiveness of integrating MoE with cutting-edge models across a diverse range of tasks [Shazeer et al., 2017, Fedus et al., 2022, Zhou et al., 2023]. These works have also tackled key challenges such as accuracy and training instability [Nie et al., 2021, Zhou et al., 2022, Puigcerver et al., 2023]. Given its ability to assign input partitions to specialized experts, MoE naturally lends itself to multimodal applications. This approach has been explored in fields such as vision-language modeling [Mustafa et al., 2022, Shen et al., 2023] and dynamic image fusion [Cao et al., 2023]. However, the application of MoE in complex real-world settings, such as those involving FlexiModal Data, remains largely unexplored. This gap presents an opportunity to leverage MoE’s potential in handling its intricate and multifaceted nature such as multimodal EHR, where reliable multimodal integration is crucial.
@@
While MoE has been widely employed to scale up large models, its theoretical foundations have remained nascent. Recently, Nguyen et al. [2023] provided convergence rates for both density and parameter estimation of Softmax gating Gaussian MoE. They connected these rates to the solvability of systems of polynomial equations under Voronoi-based loss functions. Later, Nguyen et al. [2024c] extended these theories to the top-K sparse softmax gating Gaussian MoE. Their theories further characterize the effect of the sparsity of gating functions on the behaviors of parameter estimation and verify the benefits of using top-1 sparse softmax gating MoE in practice. Other theoretical results for MoE include estimation rates of parameters and experts for multinomial logistic MoE Nguyen et al. [2024b], for dense-to-sparse gating MoE Nguyen et al. [2024a], for Gaussian gating MoE [Nguyen et al., 2024d], and for input-independent gating MoE Ho et al. [2022].
@@ B.1 MIMIC-IV and MIMIC-III Datasets
In the ICU, where rapid and informed decisions are crucial, accurate mortality prediction is essential to provide clinicians with advanced warnings of patient deterioration, aiding in critical decision-making processes Awad et al. [2017]. Similarly, the prediction of patient length-of-stay is indispensable for optimizing treatment plans, resource allocation, and discharge processes Bertsimas et al. [2022]. Further, phenotyping of critical care conditions is highly relevant to comorbidity detection and risk adjustment and presents a more challenging task than binary classification, due to the heterogeneous presentation of conditions and the larger number of prediction tasks Zhang et al. [2023]. We concentrate on three critical care tasks as highlighted in [Keuning et al., 2020], performing extensive empirical analysis on each building block of the proposed framework.
@@
  • 25-PHE  In this multilabel classification problem, we attempt to predict one of 25 acute care conditions Elixhauser [2009], Lovaasen and Schwerdtfeger [2012] (e.g., congestive heart failure, pneumonia, shock, etc.) at the end each each patient’s ICU stay. Because the original task was designed for diagnoses based on ICD-9 codes, but MIMIC-IV includes both ICD-9 and ICD-10 codes, we map patients with diagnoses coded using ICD-10 using the conversion database provided by Butler [2007].
@@
We implement an in-hospital mortality prediction (48-IHM) task to evaluate our method’s ability to predict short-term patient deterioration. Similarly, an accurate determination of patient discharge times is crucial for optimizing patient outcomes and hospital resource allocation Bertsimas et al. [2022], which motivates our length-of-stay (LOS) task. We frame 48-IHM and LOS as binary classification problems and use a 48-hour observation window (for patients who spent at least 48 hours in the ICU) to predict in-hospital mortality (48-IHM) and discharge (without expiration) within the 48 hours following the observation window (LOS). Lastly, identifying the presence of specific acute care conditions in patient records is essential for various clinical objectives, including the construction of cohorts for clinical studies and the detection of comorbidities Agarwal et al. [2016]. Traditional methods, often reliant on manual chart reviews or simple billing code-based definitions, are increasingly being supplemented by machine learning techniques Harutyunyan et al. [2019]; automating this process requires high-fidelity classifications, motivating our 25-type phenotype classification (25-PHE) task. In this multilabel classification problem, we attempt to predict one of 25 acute care conditions using data from the entire ICU stay.
@@
In our initial analysis, we focused on patients with no missing modalities, resulting in a dataset comprised of 8,770 ICU stays for the 48-IHM and LOS tasks, and 14,541 stays for the 25-PHE task. For our analyses with missing observations, we include a total of 35,129 stays for 48-IHM and LOS, and 71.173 for 25-PHE. To evaluate the single-label tasks, 48-IHM and LOS, we employ the F1-score and AUROC as our primary metrics. In line with previous studies [Zhang et al., 2023, Lin et al., 2019, Arbabi et al., 2019], we use macro-averaged F1-score and AUROC to assess the 25-PHE task.
@@
We leveraged data from MIMIC-IV [Johnson et al., 2020], a comprehensive database with records from nearly $300k$ patients admitted to a medical center from 2008 to 2019, focusing on the subset of 73,181 ICU stays. We were able to link core ICU records (containing lab results and vital signs) to corresponding chest X-rays [Johnson et al., 2019b], radiological notes [Johnson et al., 2023], and electrocardiogram (ECG) data [Gow et al., 2022] taking place during a given ICU stay. We allocated 70 percent of the data for model training, with the remaining 30 percent evenly split between validation and testing.
@@ B.2 MOSI and MOSEI Datasets
Following previous work such as Han et al. [2021b], we adopt mean absolute error (MAE), Pearson correlation (Corr), binary classification accuracy, F1 score computed for non-negative/negative class as evaluation metrics.
@@ B.4 CIFAR-10 Dataset
CIFAR-10 Krizhevsky et al. [2009] is an established computer-vision dataset used for object recognition. It consists of 60,000 32x32 color images containing one of 10 object classes (”plane”, ”car”, ”bird”, ”cat”, ”deer”, ”dog”, ”frog”, ”horse”, ”ship”, ”truck”), with 6000 images per class.
Table 7: Dataset Summary
  Dataset Research Area Modalities Sample Size Tasks
MIMIC-III Healthcare Time-Series, Text 36,212 Mortality, length-of-stay, phenotyping
MIMIC-IV Healthcare Time-Series, Text, Images, ECG 73,173 Mortality, length-of-stay, phenotyping
CMU-MOSI Affective Computing Text, Video, Audio 2,199 Sentiment
CMU-MOSEI Affective Computing Text, Video, Audio 22,777 Sentiment, emotions
PAM Healthcare Time-Series 5,333 Activity recognition
CIFAR-10 Multimedia Images 60,000 Image classification
@@ D.1 MIMIC-IV
In the preprocessing stage, we focused on 30 pertinent lab and chart events from each patient’s ICU record for vital sign measurements. For chest X-rays, we utilized a pre-trained DenseNet-121 model [Cohen et al., 2022], which was fine-tuned on the CheXpert dataset [Irvin et al., 2019], to extract 1024-dimensional image embeddings. For radiological notes, we obtained 768-dimensional embeddings using the BioClinicalBERT model [Alsentzer et al., 2019]. ECG signals were processed using a convolutional autoencoder, adapted from Attia et al. [2019], to generate a 256-dimensional embedding for each ECG.
@@
We selected 30 time series events for inclusion, following Soenksen et al. [2022]. Nine of these were vital signs: heart rate, mean/systolic/diastolic blood pressure, respiratory rate, oxygen saturation, and Glascow Coma Scale (GCS) verbal, eye, and motor response. We also included 21 lab values: potassium, sodium, chloride, creatinine, urea nitrogram, bicarbonate, anion gap, hemoglobin, hematocrit, magnesium, platlet count, phosphate, white blood cell count, total calcium, MCH, red blood cell count, MCHC, MCV, RDW, platlet count, neutrophil count, and vancomycin. We standard scale each time series value to have mean 00 and standard deviation $1$, based on the values in the training set.
@@
To incorporate a medical imaging modality into our analyses, we use the MIMIC-CXR-JPG Johnson et al. [2019a] module available from Physionet [Goldberger et al., 2000], which includes 377,110 JPG format images derived from the DICOM-based MIMIC-CXR database Johnson et al. [2019b]. Following Soenksen et al. [2022], for each image, we resize each JPG image to 224 $\times$ 224 pixels and then extract embeddings from the last layer of the Densenet121 model. We identify X-rays taken while the patient was in the ICU by first matching subject IDs in MIMIC-CXR-JPG with the core MIMIC-IV database, then limiting these matched X-rays to those with a chart time occuring between an ICU admission and discharge.
@@
To incorporate text data, we use the MIMIC-IV-Note module Johnson et al. [2023], which contains 2,321,355 deidentified radiology reports for 237,427 patients that can be matched with patients in the main MIMIC-IV via a similar approach to chest X-rays. We note that we were unable to obtain intermediate clinical notes (i.e., notes made by clinicians throughout a patient stay), as those have not yet been publicly released. We extract note embeddings using Bio-Clinical BERT [Alsentzer et al., 2019].
@@
To include ECGs as an additional modality in our models, we utilize the MIMIC-IV-ECG Gow et al. [2022] module, which includes approximately 800,000 ECGs (10 seconds, sampled at 500 Hz) collected from nearly 160,000 unique patients. To transform the ECGs so that they are suitable for input to our model, we adopt a convolutional autoencoder approach, adapted from Attia et al. [2019], that compresses each ECG into a 256-dimensional vector. Specifically, each diagnostic ECG contains a $5000\times 12$ dimensional vector (5000 time points $\times$ 12 ECG leads). To prepare the ECG for input to the autoencoder, we only include the first $4096$ time points. We then train the autoencoder to compress the ECG into a 256-dimensional latent vector, and then reconstruct the original ECG using upsampling layers, using mean squared error as our loss function. The architecture is shown in Figure 6. We train the autoencoder with 90% of the ECGs available in the MIMIC-IV-ECG projection and use the rest for validation. We selected a batch size of 2048, and reduced the learning rate by a factor of $0.5$ if the validation loss had plateaued for $3$ epochs. Training stopped if the validation loss had not decreased for $6$ epochs. For our encoder, we use filter numbers of $[16,16,32,32,64,64]$, kernel widths of $[5,5,5,3,3,3]$ and a dropout rate of $0.1$. For the decoder, we use the same filter numbers and kernel widths in reverse, and maintain a dropout rate of $0.1$.